Please solve.
3x = y + 5
x + 2y = 4
My answer is y=1, x=2 (apparently I solve differently than my son's teacher)

Found 2 solutions by rfer, solver91311:
Answer by rfer(16322)

3x=y+5 same as y=3x-5
x+2y=4
x+2(3x-5)=4
x+6x-10=4
7x=14
x=2
--------------
2+2y=4
2y=2
y=2/2
y=1
-------------
3*2=5+1
I get the same as you do.
Bob

Answer by solver91311(24713)



Can't say what the teacher is doing wrong, if anything. and is correct, and given error free arithmetic should be the result of any solution method: Substitution, Elimination, Gauss-Jordan Row Reduction, and/or Cramer's Rule. Which is the case; I checked.

The problem could be that the teacher has specified a way of presenting the solution. I.e. s/he may want the answer in ordered pair form or perhaps even in solution set form:

Ordered pair:



Solution set



Then again, it wouldn't be the first time that a teacher, textbook, and/or answer key was wrong. Won't be the last time either.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism